for a certain day the depth of water,h, in metres in Tofino,B.C at time t, in hours, is given by the formula:
h(t)= 7.8+sin[pi/6(t-3)], tE[0,24], Assume t=0 represents midnight. Provide an algebraic solution to determine the time(s) of day, the water reaches the depth of 10.29m, (Give the answer in hours and minutes)
Thank you

Answer by addingup(3677)

10.29(0)= 7.8+sin[pi/6(0-3)] = 7.8 and .8*60 = 48 so the water reaches 10.29m in 7 hours and 48 minutes